# Chip-Scale Point-Source Sagnac Interferometer by Phase-Space Squeezing

### v2

@@
Matter-wave interferometry is essential to both science and technology. Phase-space squeezing has been shown to be an advantageous source of atoms, whereby the spread in momentum is decreased. Here, we show that the opposite squeezing may be just as advantageous. As an exemplification, we analyze the effect of such a source on point source atom interferometry (PSI), which enables rotation sensing. We describe how a squeezed PSI (SPSI) increases the sensitivity and dynamic range while facilitating short cycle times and high repetition rates. We present regions in parameter space for which the figures of merit are improved by orders of magnitude and show that under some definition of compactness, the SPSI is superior by more than four orders of magnitude. The SPSI thus enables either enhancing the performance for standard size devices or maintaining the performance while miniaturizing to a chip-scale device, opening the door to real-life applications.
@@ Introduction
Since the pioneering experiments of atom interferometry (?, ?, ?, ?, ?), researchers have explored the measurement of rotation using the Sagnac effect in closed-area atom interferometers (?). These investigations have revealed exceptionally high sensitivity, comparable to state-of-the-art optical Sagnac interferometers (?). Since the early 2000s, these advancements have been driven by potential applications in inertial guidance  (?, ?), geophysics  (?) and space-based research  (?, ?, ?, ?). The majority of matter-wave interferometers studied since then rely on two-photon Raman transitions for manipulating atomic wave packets  (?, ?, ?, ?, ?, ?, ?).
Here we propose a squeezed point-source interferometer (SPSI) for increasing the dynamic range of the interferometer and its sensitivity without increasing the time of operation or, alternatively, achieving the standard sensitivity with a smaller device. Moreso, a longstanding sought-after goal is to miniaturize rotation sensing  (?, ?, ?, ?), and we show that the SPSI opens the door to such miniaturization, and even enables a chip-scale device.
@@
Our method is based on adding a stage of pre-acceleration where the initial cloud goes through phase-space squeezing. Phase-space squeezing has been discussed extensively regarding delta-kick cooling  (?, ?, ?, ?, ?), in which the spread in momentum is decreased at the expense of increasing the spread in position. In contrast, here, we propose to increase the spread in momentum while the spread in position is effectively reduced. As we show, integrating an inhomogeneous repulsive force to accelerate atom motion, either prior to or during the interferometer sequence, has the capability to substantially enhance the operational efficiency and sensitivity of a PSI device. We show that the figures of merit are enhanced by orders of magnitude, and under a definition of compactness, the enhancement in performance is about four orders of magnitude.
@@
An example of the effect of phase-space squeezing on the interferometer is illustrated in Fig. 1. The squeezing leads to both an increase in the number of oscillations within the cloud size and an increase in the contrast.
@@
Figure 1: Principle of the squeezing effect in the interferometer. The atomic phase-space distribution is shown for the standard PSI interferometer (blue) and the SPSI interferometer (orange), at two different times: (A) at the start of the interferometer sequence, $t_{\rm{acc}}$, occurring just after the acceleration stage due to the repulsive-potential pulse, and (B) at the imaging time, $t_{\text{final}}$, following the complete interferometer sequence. The phase-space distribution at $t_{\rm{final}}$ becomes tilted due to expansion, and the velocity-dependent interferometer phase introduces an oscillatory phase-space density upon detecting a single state. As shown at the bottom of (B), the latter oscillation produces a spatial density oscillation, which, upon being projected onto the position axis, forms the observed signal (output) of the interferometer. In the SPSI, when the initial phase-space distribution is squeezed with increased velocity, the subsequent expansion results in a phase-space distribution having a high aspect ratio. Upon projection onto the position axis, the SPSI signal shows improvement with more oscillations and an improved contrast.
@@ Results
An established method for constructing a Laser-Pulse Atom Interferometry (LPAI) system involves a splitting and recombining laser giving rise to stimulated transitions (e.g., Raman, Bragg, or single photons for clock transitions) to implement a $\pi/2-\pi-\pi/2$ pulse sequence. The resulting phase shift (?), for free-falling atoms, is typically described by
@@
where $\mathbf{K}_{\text{eff}}$ is the wave vector of the momentum given to the atom during the laser pulse, $\mathbf{a}$ is the constant acceleration of the platform relative to the atoms (including gravity), $T_{R}$ is the time interval between the interferometer pulses, $\mathbf{\Omega}$ represents the rotation rate of the platform, and $\mathbf{v}$ denotes the initial velocity of the atom. The second term in this equation is identified as the rotation phase. To effectively implement LPAI gyroscopes based on the $\pi/2-\pi-\pi/2$ pulse sequence, it is important to differentiate phase shifts induced by rotations from those caused by accelerations of the experimental apparatus. This can be achieved by combining the signals from two interferometers with counter-propagating atoms (?, ?, ?, ?).
@@
The rotation phase in Eq. 1, can be expressed as the Sagnac phase (?) $\Delta\Phi_{\Omega}=(2m/\hbar)\mathbf{\Omega}\cdot\mathbf{A}$. In this context, $m$ is the atomic mass, while $\mathbf{A}$ is the area enclosed by the interferometer arms. Notably, these two expressions for the phase shift are equivalent, as demonstrated through the relation $\mathbf{A}=\left(\mathbf{v}_{r}T_{R}\right)\times\left(\mathbf{v}T_{R}\right)$, where $\mathbf{v}_{r}=\hbar\mathbf{K}_{\text{eff}}/m$ is the atom’s recoil velocity. Thus, the rotation measurement in LPAI gyroscopes depends on two distinct velocities: the recoil velocity and the initial velocity of the atom. The recoil velocity is linked to the momentum imparted on the atoms by the splitting and recombining laser. In the conventional Raman method, this momentum is equivalent to $2\hbar\mathbf{K}=\hbar\mathbf{K_{\text{eff}}}$, where $\mathbf{K}$ is the wave-vector of the Raman laser, although advanced techniques have demonstrated the potential to achieve notable higher values  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). The position and velocity distribution of the atoms prior to and during the interferometer sequence is the topic of this work.
The point-source interferometer (PSI) is a specific kind of light-pulse atom interferometer in an expanding cloud of cold atoms  (?, ?, ?), and it is used in the following as a case study to examine in detail how the performance-enhancing phase-space squeezing works. The PSI method employs a single cloud of cold atoms and measures the rotation of the platform by probing the spatial frequency of the atomic density of a given output internal atomic state. The method exploits the correlation between position and velocity generated by the expansion of the cloud throughout the interferometer sequence. When the final cloud becomes much larger than the initial cloud, each atom’s final position, $\mathbf{x}$, is essentially dictated by its initial thermal velocity, $\mathbf{v}$. This allows for the approximation $\mathbf{x}\approx\mathbf{v}T_{\text{ex}}$, where $T_{\text{ex}}$ represents the total expansion time. In this limit, the Sagnac phase becomes  (?)
$\Delta\Phi_{\Omega,\text{PS}}=\frac{2{T_{R}}^{2}}{T_{\text{ex}}}\left(\mathbf{K}_{\text{eff}}\times\mathbf{\Omega}\right)\cdot\mathbf{x}\equiv\mathbf{k}_{x}\cdot\mathbf{x}.$ (2)
This phase is subsequently encoded as a spatial fringe pattern onto the image of the expanded cloud of atoms at the conclusion of the interferometer sequence. It is important to highlight that any deviation from the established position-velocity correlation results in a reduction of contrast within the fringe pattern, consequently diminishing the precision of the rotation measurement. Thus, to attain an approximation of a point source, it is necessary to cool and confine the atoms, resulting in reduced thermal velocities. However, this in turn leads to a smaller interferometer area, thus yielding lower resolution. Therefore, to enhance the resolution of the rotation phase measurement, it is advantageous to increase these velocities while preserving the position-velocity correlation.
@@
According to Eq. 2, the probability of an atom with initial velocity $\mathbf{v}$ to be at the end of the interferometric sequence in a given internal state $|s\rangle$ is $P_{s}({\bf v})=1/2\cdot[1+\cos(\mathbf{k}_{v}\cdot\mathbf{v}+\phi_{s})]$, where $\mathbf{k}_{v}=2(\mathbf{K}_{\rm eff}\times\mathbf{\Omega})T_{R}^{2}$ and $\phi_{s}$ is determined by gravity and platform acceleration. In an ideal case, the atomic cloud starts to expand freely from a point source, and the position of each atom is linearly proportional to its velocity $\mathbf{x}=\mathbf{v}T_{\rm{ex}}$, where $T_{\rm{ex}}$ is the time of expansion. In this case, the atoms form a spatial density that oscillates with a wave-vector $\mathbf{k}_{x}$, according to Eq. 2, from which one extracts the angular velocity component, which is transverse to the plane defined by the vectors $K_{\rm{eff}}$ and $k_{x}$.
@@
In the more realistic case, we may consider an initial cloud with an isotropic Gaussian spatial distribution having a width $\sigma_{x0}$ and a Gaussian velocity distribution having a width $\sigma_{v0}$. In this case, one can show that the output density has the form  (?)
$\rho({\bf x})\propto e^{-x^{2}/2\sigma_{f}^{2}}[1+C\cos(\alpha\mathbf{k}_{x}\cdot{\bf x}+\phi_{s})],$ (3)
where $\sigma_{f}=\sqrt{\sigma_{x0}^{2}+T_{\rm{ex}}^{2}\sigma_{v0}^{2}}$ is the final cloud width, $\alpha=1-\sigma_{x0}^{2}/\sigma_{f}^{2}$, and the contrast $C$ is given by
$C=\exp\left[-\frac{1}{2}k_{x}^{2}\sigma_{x0}^{2}\left(1-\frac{\sigma_{x0}^{2}}{\sigma_{f}^{2}}\right)\right]\,.$ (4)
@@
The sensitivity of an interferometer to rotation, limited by projection noise, hinges on two additional factors: the number of atoms per operation ($N$) and the frequency of operations ($\nu$). This sensitivity ($\delta\Omega$) is given by the expression  (?):
$\delta\Omega=\frac{T_{\rm{ex}}/T_{R}^{2}}{2C\sqrt{N\nu/2}K_{\text{eff}}\sigma_{f}}\,,$ (5)
@@
where it is preferable to use a longer interferometer duration ($T_{R}$) and enlarge the cloud size ($\sigma_{f}$) to achieve good sensitivity.
@@
In contrast to PSI, in interferometric rotation sensors, where the signal is atom population or light intensity in two output ports, the measured quantity is the Sagnac phase, which typically exhibits a linear correlation with the rotation. The smallest measurable phase difference determines the smallest detectable change in rotation, whether from zero rotation or from any given rotation. Hence, sensitivity characterizes both the smallest detectable rotation and the uncertainty in detecting a given rate of rotation. On the other hand, in the PSI, the minimum detectable rotation by the interferometer is determined by the condition that the cloud size exceeds the period of spatial oscillation, expressed as $k_{x}\sigma_{f}>1$. In addition, the contrast diminishes if the rotation rate exceeds the threshold where the fringe period becomes larger than the initial cloud size, namely $k_{x}\sigma_{x0}>1$. This implies a constraint on the fringe periodicity within the range $1/\sigma_{f}<k_{x}<1/\sigma_{x0}$, leading to a limitation on the dynamic range of the interferometer, namely, $\Omega_{\rm min}<\Omega<\Omega_{\rm max}$, where from Eq. 2 (and neglecting $\alpha$ from Eq. 3) we find
@@
such that $\Omega_{\rm max}/\Omega_{\rm min}\sim\sigma_{f}/\sigma_{x0}$. These limits are approximations closely related to the method used for image analysis. Previous works discussed this issue within various methods, including the ellipse fitting procedure  (?), phase shear  (?), or phase map  (?). Finally, it should be noted that by combining Eqs. 5 and 6, the sensitivity in detecting a rotation rate, $\delta\Omega$, is associated with the minimally detectable rate, $\Omega_{\rm{min}}$, by the relationship $\delta\Omega\approx 1/\sqrt{N/2}$. This suggests that for PSI, the sensitivity ($\delta\Omega$) solely pertains to uncertainty and may be much smaller (better) than the smallest detectable rotation rate.
Figure 2: Contrast and dynamic range. (A) Contrast vs. angular velocity ($\Omega$). Using 87Rb atoms, the simulation parameters are initial cloud size $\sigma_{x0}=100\,\rm{\mu m}$, temperature $T=5\,\rm{\mu K}$, and time between pulses $T_{R}=5\,\rm ms$. For the repulsive potential, we take a beam power of $P=1\,$W with a cross-section of $400\crossproduct 400\,\mu$m2, blue-detuned by $\Delta=2\pi\cdot 10\,$GHz, and focused such that in the direction of acceleration it gives rise to an inverted harmonic potential and in the transverse direction a constant potential. The harmonic profile was optimized for achieving a large squeezing factor before the atoms exceed the region where the repulsive potential is harmonic. Following Eqs. 8 and 11, and choosing $t_{\text{acc}}=80.6\,\mu$s, we find a squeezing parameter of 22.81. The purple line represents the analytical solution of PSI contrast, while the dashed orange line corresponds to SPSI, according to Eq. 4. The data points are the result of a numerical simulation. While the contrast reduction due to the ratio between the fringe periodicity and the initial cloud size determines the upper limit of the detection range, there exists a lower limit when the fringe period becomes larger than the final cloud size. These limits are roughly given by $\Omega_{\rm min}$ and $\Omega_{\rm max}$ (Eq. 6), which are presented in the graph as vertical dashed lines. It is evident that using the SPSI improves the contrast greatly, thereby increasing the detection dynamic range. At high angular velocities, the contrast of the data points decreases faster than the analytical solution due to the short spacing between fringes relative to the detection pixel size considered only in the numerical calculation. (B) The analytical curve of the detection range as a function of $\eta$ according to Eq. 6. The dashed line presents the SPSI simulated in (A) with $\eta=22.81$.
@@
Integrating a repulsive potential that varies spatially to accelerate atom motion, either prior to or during the interferometer sequence, has the capability to substantially enhance the operational efficiency and sensitivity of a PSI device. Repulsive forces have been discussed in numerous contexts  (?, ?, ?, ?, ?, ?), but here, the repulsive potential has the role of enlarging the area enclosed by the interferometer arms, while maintaining and even improving the crucial position-velocity correlation. This is achieved by applying a repulsive potential for an acceleration time $t_{\rm acc}$. Let us define the coordinate system such that the $z$ axis is along the direction of the splitting and recombining laser beam. We may apply a quadratic repulsive potential along one or two axes transverse to $z$, but for the sake of simplicity, we describe only the dynamics along the $x$ coordinate. If the repulsive potential is quadratic $V_{\rm r}(x)=-(1/2)m\omega^{2}x^{2}$, where $m$ is the atomic mass, then after the acceleration stage, the initial coordinates in the direction $x$ of an atom are transformed as
$x(t_{\rm acc})=c\cdot x(0)+s\cdot v_{x}(0)/\omega\,,\quad v_{x}(t_{\rm acc})=\omega s\cdot x(0)+c\cdot v_{x}(0)\,,$ (7)
where $c=\cosh(\omega t_{\rm acc})$ and $s=\sinh(\omega t_{\rm acc})$.
It is easy to show that the phase-space distribution that forms after the repulsive pulse is exactly equivalent to the distribution that forms after free propagation for an effective duration $t_{\rm eff}$ when the initial distribution is squeezed with effective position uncertainty $\tilde{\sigma}_{x0}=\sigma_{x0}/\eta$ and velocity uncertainty $\tilde{\sigma}_{v0}=\eta\sigma_{v0}$, with the squeezing parameter being
$\eta\equiv\sqrt{c^{2}+\frac{\omega^{2}\sigma_{x0}^{2}}{\sigma_{v0}^{2}}s^{2}},$ (8)
and the effective time is given by
$t_{\rm eff}=\frac{sc}{\omega\eta^{2}}\left(1+\frac{\omega^{2}\sigma_{x0}^{2}}{\sigma_{v0}^{2}}\right)\,.$ (9)

### v3

@@
Matter-wave interferometry plays a significant role in scientific research and technological applications. While position-momentum phase-space squeezing has been demonstrated to increase the coherence of atom sources by reducing momentum spread, we theoretically investigate the potential advantages of the opposite squeezing. As a case study, we analytically and numerically examine its effect on point source atom interferometry (PSI) for rotation sensing. Our analysis reveals that this squeezed PSI (SPSI) approach can significantly improve sensitivity and dynamic range while enabling shorter cycle times and higher repetition rates. Through simulations, we identify parameter spaces where sensitivity and dynamic range are enhanced by orders of magnitude. Under a specific definition of compactness, our calculations show that SPSI outperforms standard PSI by over four orders of magnitude. These theoretical findings suggest that SPSI could either enhance performance in standard-sized devices or maintain performance in miniaturized chip-scale devices, potentially paving the way for new practical applications.
@@ Introduction
Since the pioneering experiments of atom interferometry (?, ?, ?, ?, ?), researchers have explored the measurement of rotation using the Sagnac effect in closed-area atom interferometers (?). These investigations have revealed exceptionally high sensitivity, comparable to state-of-the-art optical Sagnac interferometers (?). Since the early 2000s, these advancements have been driven by potential applications in inertial guidance  (?, ?), geophysics  (?) and space-based research  (?, ?, ?, ?). The majority of matter-wave interferometers studied since then are laser pulse atom interferometers (LPAIs) that rely on two-photon Raman transitions for splitting, diverting, and recombining the atomic wave packets  (?, ?, ?, ?, ?, ?, ?).
The sensitivity of LPAIs to linear acceleration relies on a space-time area between two interferometer arms, while sensitivity to rotations requires that the two arms enclose a real space area. To effectively implement LPAI gyroscopes based on the $\pi/2-\pi-\pi/2$ pulse sequence, it is important to differentiate phase shifts induced by rotations from those caused by accelerations of the carrying platform. This can be achieved by combining the signals from two interferometers with counter-propagating atoms  (?, ?, ?, ?). The point-source interferometer (PSI) is a specific kind of light-pulse atom interferometer  (?, ?, ?) that allows sensing both acceleration and rotation by probing the phase and spatial frequency of a fringe pattern formed by a given internal atomic state in an expanding cloud of cold atoms. The method exploits the correlation between position and velocity generated by the expansion of the cloud throughout the interferometer sequence.
The compact nature of the PSI setup makes it suitable for applications where space is limited, such as in portable devices or space-based research and it has the potential to achieve the long sought-after goal of miniaturizing rotation sensing  (?, ?, ?, ?). However, the sensitivity and dynamic range of PSI is limited because on the one hand the initial cloud should be as cold as possible to allow it to be confined to a small volume but on the other hand it should have a fast expansion rate in order to quickly achieve a good position-velocity correlation and develop a highly resolvable spatial fringe pattern.
Here, we propose a squeezed point-source interferometer (SPSI) for increasing the dynamic range of the interferometer and its sensitivity or, alternatively, achieving the standard sensitivity with a smaller device, or even enable a chip-scale sensor. Our method is based on adding a stage of pre-acceleration where the initial cloud undergoes position-momentum phase-space squeezing. Phase-space squeezing has been discussed extensively regarding delta-kick cooling  (?, ?, ?, ?, ?), in which the spread in momentum is decreased at the expense of increasing the spread in position. In contrast, here, we propose to increase the spread in momentum while the projected spread in position is reduced, in the sense that if the atomic dynamics is reversed, the atoms concentrate in a volume smaller than their actual initial volume.
Our proposed SPSI sequence consists of several key stages. Initially, a thermal cloud of cold atoms is prepared using standard cooling techniques. This is followed by a novel pre-acceleration phase where the atoms experience a repulsive potential, inducing position-momentum phase-space squeezing. Subsequently, the atoms undergo a conventional three-pulse interferometry sequence $(\pi/2-\pi-\pi/2)$. Finally, the atomic state is detected, revealing a spatial fringe pattern that encodes rotation information. While our current study focuses on theoretical analysis and numerical simulations of this sequence, it provides a foundation for future experimental implementations.
@@
Through our analytical and numerical studies, we demonstrate that integrating an inhomogeneous repulsive force to accelerate atom motion, prior to the interferometer sequence, has the capability to substantially enhance the operational efficiency and sensitivity of a PSI device. Our calculations show that the figures of merit are enhanced by orders of magnitude, and under a definition of compactness, the enhancement in performance is about four orders of magnitude.
@@
An example of the effect of phase-space squeezing on the interferometer is illustrated in Fig. 1. Our simulations indicate that the squeezing leads to both an increase in the number of oscillations within the cloud size and an increase in the contrast.
@@
Figure 1: Principle of the squeezing effect in the interferometer. The atomic phase-space distribution is shown for the standard PSI interferometer (blue) and the squeezed-PSI (SPSI) interferometer (orange), at two different times: (A) at the start of the interferometer sequence, $t_{\rm{acc}}$, occurring just after the acceleration stage due to the repulsive-potential pulse, and (B) at the imaging time, $t_{\text{final}}$, following the complete interferometer sequence. The phase-space distribution at $t_{\rm{final}}$ becomes tilted due to expansion, and the velocity-dependent interferometer phase introduces an oscillatory phase $\Delta\phi(v_{x})=2k_{\rm eff}\Omega_{y}T_{R}^{2}v_{x}\equiv k_{v}v_{x}$ that modulates the phase-space distribution as $\rho(x,v_{x})\to\frac{1}{2}\rho(x,v_{x})[1+\cos(k_{v}v_{x})]$ upon detecting a single state. As shown at the bottom of (B), the latter oscillation produces a spatial density oscillation, which, upon being projected onto the position axis, forms the observed signal (output) of the interferometer. In the SPSI, when the initial phase-space distribution is squeezed with increased velocity, the subsequent expansion results in a phase-space distribution having a high aspect ratio. Upon projection onto the position axis, the SPSI signal shows improvement with more oscillations and an improved contrast. This figure was generated using a simplified version of the numerical simulations described in the Materials and Methods section, focusing on illustrating the key principles of phase-space evolution in both PSI and SPSI configurations.
@@ Results
An established method for constructing a Laser-Pulse Atom Interferometry (LPAI) system involves a splitting and recombining laser giving rise to stimulated transitions (e.g., Raman, Bragg, or single photons for clock transitions) to implement a $\pi/2-\pi-\pi/2$ pulse sequence. In an interferometer using freely falling atoms, the interferometer phase shift due to acceleration and rotation of the platform is given by  (?, ?)
@@
where $\mathbf{K}_{\text{eff}}$ is the wave vector of the momentum given to the atom during the laser pulse, $\mathbf{a}$ is the acceleration of the platform relative to the atoms (including gravity), which is assumed to be constant during the measurement, $T_{R}$ is the time interval between the interferometer pulses, $\mathbf{\Omega}$ represents the rotation rate of the platform, and $\mathbf{v}$ denotes the initial velocity of the atom. The second term in this equation is identified as the rotation phase.
@@
The rotation phase in Eq. 1, can be expressed as the Sagnac phase (?) $\Delta\Phi_{\Omega}=(2m/\hbar)\mathbf{\Omega}\cdot\mathbf{A}$. In this context, $m$ is the atomic mass, while $\mathbf{A}$ is the area enclosed by the interferometer arms. Notably, these two expressions for the phase shift are equivalent, as demonstrated through the relation $\mathbf{A}=\left(\mathbf{v}_{r}T_{R}\right)\times\left(\mathbf{v}T_{R}\right)$, where $\mathbf{v}_{r}=\hbar\mathbf{K}_{\text{eff}}/m$ is the atom’s recoil velocity. In the conventional scheme, the momentum transfer in each Raman pulse is equal to the difference of the wave vectors of the two counter-propagating Raman beams, such that $\mathbf{K}_{\text{eff}}\approx 2\mathbf{K}$, where $\mathbf{K}$ is the wave vector of a single Raman laser. Advanced techniques have demonstrated considerably higher values of $\mathbf{K}_{\text{eff}}$  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?).
In the following we use the point-source interferometer (PSI) as a case study to examine in detail how position-momentum phase-space squeezing enhances the performance of the interferometer. The position and velocity distribution of the atoms prior to and during the interferometer sequence and its effect on performance is the topic of this work. The PSI exploits the correlation between position and velocity generated by the expansion of the cloud throughout the interferometer sequence. According to Eq. 1, the probability that an atom with initial velocity $\mathbf{v}$ is at the end of the interferometric sequence in a given internal state $|s\rangle$ is
$P_{s}({\bf v})=1/2\cdot[1+\cos(\mathbf{k}_{v}\cdot\mathbf{v}+\phi_{s})],$ (2)
@@
where $\mathbf{k}_{v}=2(\mathbf{K}_{\rm eff}\times\mathbf{\Omega})T_{R}^{2}$ and $\phi_{s}$ is determined by gravity and acceleration of the platform. In an ideal case, the atomic cloud starts to expand freely from a point source, and the position of each atom is linearly proportional to its velocity $\mathbf{x}=\mathbf{v}T_{\rm{ex}}$, where $T_{\rm{ex}}$ is the time of expansion. In this limit, the Sagnac phase becomes  (?)
$\Delta\Phi_{\Omega,\text{PS}}(\mathbf{x})=\frac{2{T_{R}}^{2}}{T_{\text{ex}}}\left(\mathbf{K}_{\text{eff}}\times\mathbf{\Omega}\right)\cdot\mathbf{x}\equiv\mathbf{k}_{x}\cdot\mathbf{x}.$ (3)
In this case, the atoms form a spatial density constituting a fringe pattern that oscillates with a wave-vector $\mathbf{k}_{x}$, from which one extracts the angular velocity component, transverse to the plane defined by the vectors $\mathbf{K}_{\text{eff}}$ and $\mathbf{k}_{x}$. It is important to highlight that any deviation from the established position-velocity correlation results in a reduction of contrast within the fringe pattern, consequently diminishing the precision of the rotation measurement. Thus, to attain an approximation of a point source, it is necessary to cool and confine the atoms, resulting in reduced thermal velocities. However, this, in turn, leads to a smaller interferometer area, thus yielding a lower resolution. Therefore, to enhance the resolution of the rotation phase measurement, it is advantageous to increase these velocities while preserving the position-velocity correlation.
@@
Let us now examine the more general case of an initial atomic cloud with a finite size. We consider a thermal cloud with an initial isotropic Gaussian spatial distribution having a width $\sigma_{x0}$ and a Gaussian velocity distribution having a width $\sigma_{v0}$. In this case, one can show (see supplementary information for an alternative derivation for quantum wave packets) that the output density has the form  (?, ?)
$\rho_{f}({\bf x})\propto e^{-x^{2}/2\sigma_{f}^{2}}[1+C\cos(\alpha\mathbf{k}_{x}\cdot{\bf x}+\phi_{s})],$ (4)
where $\sigma_{f}=\sqrt{\sigma_{x0}^{2}+T_{\rm{ex}}^{2}\sigma_{v0}^{2}}$ is the final cloud width, $\alpha=1-\sigma_{x0}^{2}/\sigma_{f}^{2}$, and the contrast $C$ (Eq. 5) drops considerably when the angular velocity is large such that the fringe period $2\pi/k_{x}$ becomes as small as the initial cloud size (for $\alpha\sim 1$)
$C=\exp\left[-\frac{1}{2}\alpha k_{x}^{2}\sigma_{x0}^{2}\right]\,.$ (5)
The angular velocity $\Omega$ is deduced from the measured atomic density by fitting it to the form of Eq. 4 and then extracting $\Omega$ from the periodicity $\alpha k_{x}$. This periodicity includes a deviation
$\Delta\Omega=-\frac{\sigma_{x0}^{2}}{\sigma_{f}^{2}}\Omega\,,$ (6)
which is caused by the non-ideal point-source approximation and may be eliminated to some extent if the ratio between the initial and final cloud sizes is known. The appearance of the initial cloud size $\sigma_{x0}$ in both Eqs. 5 and 6 emphasizes the importance of operation close to the point-source approximation $\sigma_{x0}\ll\sigma_{f}$, $k_{x}\sigma_{x0}\ll 1$.
@@
The sensitivity of an interferometer to rotation, limited by projection noise, hinges on two additional factors: the number of atoms at each interferometer cycle ($N$) and the number of cycles per unit time ($\nu$). As far as the deviation $\Delta\Omega$ of Eq. 6 is neglected, this sensitivity ($\delta\Omega$) is given by the expression  (?):
$\delta\Omega=\frac{\delta\Omega_{\rm single}}{\sqrt{\nu}}=\frac{T_{\rm{ex}}/T_{R}^{2}}{2C\sqrt{N\nu/2}K_{\text{eff}}\sigma_{f}}\,,$ (7)
@@
where it is preferable to use a longer interferometer duration ($T_{R}$) and enlarge the cloud final size ($\sigma_{f}$) to achieve good sensitivity.
@@
Most Sagnac interferometers that serve as rotation sensors measure the Sagnac phase directly by probing the total atomic population or light intensity at an output port of a single loop. As the angular velocity of rotation is linearly proportional to the Sagnac phase, it follows that the smallest detectable rotation change is proportional to the smallest measurable phase change. Hence, sensitivity characterizes both the smallest detectable rotation and the uncertainty in detecting a given rate of rotation. By contrast, in the PSI, the minimum detectable rotation by the interferometer is determined by the condition that the cloud size exceeds the period of spatial oscillation, expressed as $k_{x}\sigma_{f}>1$. In addition, the contrast diminishes if the rotation rate exceeds the threshold where the fringe period becomes larger than the initial cloud size, namely $k_{x}\sigma_{x0}>1$. This implies a constraint on the fringe periodicity within the range $1/\sigma_{f}<k_{x}<1/\sigma_{x0}$, leading to a limitation on the dynamic range of the interferometer, namely, $\Omega_{\rm min}<\Omega<\Omega_{\rm max}$, where from Eq. 3 (and neglecting $\alpha$ from Eq. 4) we find
@@
such that $\Omega_{\rm max}/\Omega_{\rm min}\sim\sigma_{f}/\sigma_{x0}$. The exact value of the minimally detectable rotation $\Omega_{\rm min}$ depends on the specific method used for image analysis. In this context, previous works have discussed, for example, the ellipse fitting procedure  (?), phase shear  (?), or phase map  (?). Finally, it should be noted that by combining Eqs. 7 and 8, the sensitivity in detecting a rotation rate, $\delta\Omega$, is associated with the minimally detectable rate, $\Omega_{\rm{min}}$, by the relationship $\delta\Omega/\Omega_{\rm{min}}\approx 1/\sqrt{N\nu/2}$. This suggests that for PSI, the sensitivity ($\delta\Omega$) solely pertains to uncertainty and may be much smaller (better) than the lowest detectable rotation rate.
Figure 2: Contrast and dynamic range. (A) Contrast vs. angular velocity ($\Omega$). Using 87Rb atoms, the simulation parameters are initial cloud size $\sigma_{x0}=100\,\rm{\mu m}$, temperature $T=5\,\rm{\mu K}$, and time between pulses $T_{R}=5\,\rm ms$. For the repulsive potential, we take a beam power of $P=1\,$W with a cross-section of $400\crossproduct 400\,\mu$m2 (according to the beam profile of Eq. 13), blue-detuned by $\Delta=2\pi\cdot 10\,$GHz, and focused such that in the direction of acceleration it gives rise to an inverted harmonic potential and in the transverse direction a constant potential. The harmonic profile was optimized to achieve a large squeezing factor before the atoms exceed the region where the repulsive potential is harmonic. Following Eqs. 10 and 13, and choosing $t_{\text{acc}}=80.6\,\mu$s, we find a squeezing parameter of 22.81. The purple line represents the analytical solution of PSI contrast, while the dashed orange line corresponds to SPSI, according to Eq. 5. The simulation points are the result of a numerical simulation. While the contrast reduction due to the ratio between the fringe periodicity and the initial cloud size determines the upper limit of the detection range, there exists a lower limit when the fringe period becomes larger than the final cloud size. These limits are roughly given by $\Omega_{\rm min}$ and $\Omega_{\rm max}$ (Eq. 8), which are presented in the graph as vertical dashed lines. It is evident that using the SPSI greatly improves the contrast, thereby increasing the detection dynamic range. At high angular velocities, the contrast obtained from simulation decreases faster than the analytical solution due to the short spacing between fringes relative to the detection pixel size considered only in the numerical calculation. (B) The analytical curve of the detection range as a function of $\eta$ according to Eq. 8. The dashed line presents the SPSI simulated in (A) with $\eta=22.81$.
@@
Integrating a repulsive potential that varies spatially to accelerate atom motion, prior to the interferometer sequence, has the capability to substantially enhance the operational efficiency and sensitivity of a PSI device. Repulsive forces have been discussed in numerous contexts  (?, ?, ?, ?, ?, ?), but here, the repulsive potential has the role of enlarging the area enclosed by the interferometer arms, while improving the crucial position-velocity correlation. This is achieved by applying a repulsive potential for an acceleration time $t_{\rm acc}$. To describe this position-momentum squeezing, let us define the coordinate system such that the $z$ axis is along the direction of the splitting and recombining laser beam. We may apply a quadratic repulsive potential along one or two axes transverse to $z$, but for the sake of simplicity, we describe only the dynamics along the $x$ coordinate. If the repulsive potential is quadratic $V_{\rm r}(x)=-(1/2)m\omega^{2}x^{2}$, where $m$ is the atomic mass, then after the acceleration stage, the initial coordinates in the direction $x$ of an atom are transformed as
$x(t_{\rm acc})$ $=\cosh(\omega t_{\rm acc})\cdot x(0)+\sinh(\omega t_{\rm acc})\cdot v_{x}(0)/\omega\,,$ (9)
$v_{x}(t_{\rm acc})$ $=\omega\sinh(\omega t_{\rm acc})\cdot x(0)+\cosh(\omega t_{\rm acc})\cdot v_{x}(0)\,,$
It is easy to show that the phase-space distribution that forms after the repulsive pulse is exactly equivalent to the distribution that forms after free propagation for an effective duration $t_{\rm eff}$ when the initial distribution is squeezed with effective position uncertainty $\tilde{\sigma}_{x0}=\sigma_{x0}/\eta$ and velocity uncertainty $\tilde{\sigma}_{v0}=\eta\sigma_{v0}$, with the position-momentum phase-space squeezing parameter being
$\eta\equiv\sqrt{\cosh^{2}(\omega t_{\rm acc})+\frac{\omega^{2}\sigma_{x0}^{2}}{\sigma_{v0}^{2}}\sinh^{2}(\omega t_{\rm acc})}.$ (10)
The effective time is given by
$t_{\rm eff}=\frac{\sinh(2\omega t_{\rm acc})}{2\omega\eta^{2}}\left(1+\frac{\omega^{2}\sigma_{x0}^{2}}{\sigma_{v0}^{2}}\right)\,.$ (11)
@@
In a PSI, the final size of the cloud is limited not only by the duration of expansion but also by the extent of the Raman beam along the transverse direction. The intensity profile over the cloud size must be flat to ensure that the Rabi pulses are complete and do not introduce additional phase. In our analysis, we assume that a wide beam is used to ensure this requirement, although this might involve beam sizes of centimeters or a special beam-shaping procedure. In addition, we note that the velocity along the beam axis is limited by the requirement that velocity selection of the Raman beam will not cause inefficient transitions for a large part of atoms with the largest velocities. In our SPSI procedure, the atoms are accelerated along the transverse direction and hence this will not affect their interaction with the Raman beams unless the rotation frequency during the interferometer sequence is so large that transverse velocity components will have a considerable projection along the instantaneous longitudinal direction of the Raman beams.